# Towards more realistic evaluation of LLM-based code generation: an experimental study and beyond

### v1

@@ 5.3. RQ3: How different code descriptions in the dataset affect code generation performance?
During the benchmark collection process, we notice that some projects tend to write detailed docstrings, while others prefer to describe the function’s functionality using only brief statements. To make our benchmark more diverse and reflective of real-world development scenarios, we manually write both brief and detailed docstrings for each function in HumanEvo. In this experiment, we evaluate the performance of the selected LLMs on two types of manually written docstrings. As shown in Table 6, experimental results indicate that the majority of LLMs perform better on detailed docstrings compared to brief docstrings. This is because detailed docstrings provide a more comprehensive description of the functionality of the target function, some even including the implementation logic of the target function. In addition, detailed docstrings also provide explanations for the input and output of the target function, enabling LLMs to better understand the tasks to be performed. Conversely, brief docstrings only introduce the intention of the target function, which may lead to LLMs struggling to understand user intent in certain complex scenarios, ultimately resulting in decreased performance.
@@
However, the decline in LLMs’ performance is not significant, mainly because we adopted a retrieval-augmented generation approach, retrieving a large amount of context from the project repository through similarity calculation. These contexts contribute to a lengthy prompt. However, when LLMs deal with overly long prompts, their attention gets dispersed, leading to the neglect of some useful information by the model.
RQ3 Summary: Compared to brief docstrings, detailed docstrings provide more beneficial information to LLMs, resulting in better performance on detailed docstrings. However, due to the extensive project context required for completing the repository-level code generation task, the model’s attention gets dispersed. Consequently, the performance improvement on detailed docstrings is not substantial.
Table 6. Model performance under different docstrings styles.DepSCode is short for DeepSeekCoder.
@@
6. Threads To Validity
First, due to limited computational resources, our experiments did not cover all code LLMs, thus not fully reflecting the performance of all LLMs in actual development scenarios, which may slightly affect the representativeness of our experiments. Second, currently HumanEvo only includes two languages. In the future, we will continue to expand HumanEvo to cover as many programming languages as possible. Third, in prompt format selection, we ultimately chose the one that yielded the best results after trying several formats. However, since we did not cover all available prompt formats, the one we selected may not be the best. Finally, our benchmark includes GitHub data from before 2023, which might already be present in the pretraining dataset of the models, posing a risk of data leakage. To mitigate this risk, we construct an automated benchmark construction pipeline that continuously updates the dataset.
7. Related Work
@@
CoderEval (Yu et al., 2024) contains 230 samples from 43 Python projects and 230 samples from 10 Java projects. They categorize task instances into six runnable-levels to better analyze LLMs’ code generation performance. EvoCodeBench (Li et al., 2024a) collects data from 25 real-world repositories. It also ensures that the distribution of programming tasks (with or without dependencies) matches the distribution of real repositories. RepoCoder (Zhang et al., 2023b) proposed RepoEval, which consists of the latest and high-quality real-world repositories covering scenarios such as line completion, API invocation, and function body completion. Currently, most benchmarks for LLM code generation provide simple functions, mostly targeting general functions rather than specific domains, thus limiting practicality. However, BioCoder (Tang et al., 2024), a benchmark for bioinformatics code generation, has been introduced. BioCoder carefully selects 28 high-quality projects related to bioinformatics, then parses these repositories to generate ASTs and obtain relevant data. A custom code context is written for each filtered function, including necessary imports, cross-file dependencies, and some test cases.
@@
However, all the benchmarks mentioned above have a common flaw: they all overlook the evolution of projects over time. These benchmarks take the latest version of the project repository as the context source, and such an evolving-ignored setting makes it difficult for them to accurately reflect the performance of LLMs in real-world development scenarios. Therefore, in this work, we propose HumanEvo, the first repository-level evolving-aware code generation benchmark. Before obtaining the project context for the target function, we roll back the project repository to the state before the target function was committed to ensure that the project context provided to LLMs does not suffer from issues such as future context leakage and missing useful context. Our objective is to offer LLMs a more realistic development scenario, thus more accurately reflecting the actual performance of LLMs in real-world development scenarios.
8. Conclusion
@@
In this paper, we highlight a previously overlooked flaw in evaluating existing repository-level code generation methods: the evolution of projects over time. To address it, we introduce HumanEvo, an evolving-aware repository code generation benchmark, and conduct a comprehensive empirical study to explore LLMs’ performance in pragmatic development scenarios. Experimental results demonstrate that the evolving-ignored situation leads to the inflated performance of LLMs. By incorporating temporal considerations into the benchmark, HumanEvo provides a more comprehensive assessment of model performance in pragmatic development scenarios. We believe that our empirical results and HumanEvo will offer valuable insights for the future development of LLMs-based repository-level code generation benchmarks.

### v2

@@ V-C RQ3: How different code descriptions in the dataset affect code generation performance?
During the benchmark construction process, we notice that, in real-world development, some projects tend to write detailed docstrings, while the others prefer to describe the function’s functionality using only several brief statements. Since these docstrings are used as input to models which may significantly affect model performance, we explore whether the impact of the evolution-ignored setting on model performance varies under different docstring styles. To this end, we manually write both brief and detailed docstrings for each task instance in HumanEvo. In this experiment, we evaluate the performance of the studied LLMs on the two types of manually written docstrings. As shown in Table VI, experimental results indicate that under the evolution-ignored setting, using different styles of docstrings as input consistently leads to inflated performance, and the style of the docstring does not significantly change the impact of the evolution-ignored setting on the inflated performance evaluation. Additionally, the majority of LLMs perform better on detailed docstrings compared to the brief docstrings. We suspect that the detailed docstrings provide a more comprehensive description of the functionality of the target function and include the implementation logic of the target function. In addition, detailed docstrings also provide explanations for the input and output of the target function, enabling LLMs to better understand the tasks to be addressed. Conversely, brief docstrings only introduce the intention of the target function, which may lead to LLMs struggling to understand user intent in certain complex scenarios, ultimately resulting in decreased performance.
@@
However, the decline in LLMs’ performance is not significant. We suspect that the reason could possibly be related to our use of a retrieval-augmented generation approach, by which we have retrieved a substantial amount of context from the project through similarity calculations. These contexts contribute to a lengthy prompt. However, when LLMs deal with overly long prompts, their attention gets dispersed, leading to the neglect of some useful information by the model.
RQ3 Summary: From the experimental results, we do not observe any correlation between docstring styles and evolution settings. Compared to brief docstrings, detailed docstrings provide more beneficial information to LLMs, resulting in better performance. However, the performance improvement on detailed docstrings is not substantial.
VI Threats to Validity
We have identified the following potential threats that may affect the validity of our study. Firstly, due to limited computational resources, our experiments were not able to cover all the available code LLMs, thus not fully reflecting the performance of all LLMs in actual development scenarios, which may slightly affect the representativeness of our experiments. Secondly, currently HumanEvo only includes two programming languages. In the future, we plan to address this limitation by continuously expanding HumanEvo to cover as many programming languages as possible, thereby providing a more comprehensive evaluation. Thirdly, in prompt format selection, we ultimately chose the one that yielded the best results after trying several formats. However, since we did not cover all available prompt formats, the one we selected may not be the optimal one. Lastly, our benchmark includes GitHub data before 2023, which might already be present in the pretraining dataset of the models we studied, which poses a risk of data leakage. To mitigate this risk, we have implemented an automated benchmark construction pipeline that continuously updates the dataset and rewrited the docstrings for every task instance in HumanEvo.
TABLE VI: Model performance under different docstrings styles.
@@ VI Threats to Validity
VII Related Work
VII-A Code Generation Benchmarks
Previously, the vast majority of benchmarks were constructed to evaluate the performance of LLMs in generating standalone functions [44, 45, 46, 47, 34, 26, 27, 79, 80, zhong2024memorybank], with the most representative ones being HumanEval  [19] and MBPP  [11]. HumanEval comprises 164 manually crafted programming problems covering language comprehension, reasoning, algorithms, and simple mathematics. MBPP consists of 974 short Python function problems primarily designed for novice programmers.
@@ VII-A Code Generation Benchmarks
Finding previous benchmarks fail to reflect complex programming scenarios, researchers propose CoderEval [13], the first repository-level code generation benchmark, to evaluate the performance of existing LLMs in real-world development by prompting them to generate functions selected in actual complex software. CoderEval meticulously selects 230 Python and 230 Java programming tasks from popular real-world open-source projects, evaluating the functional correctness of generated code based on execution. EvoCodeBench [64] collects data from 25 real-world repositories. It also ensures that the distribution of programming tasks (with or without dependencies) matches the distribution of real repositories. RepoCoder [6] proposed RepoEval, which consists of the latest and high-quality real-world repositories covering scenarios such as line completion, API invocation, and function body completion. Currently, most benchmarks for LLM code generation provide simple functions, mostly targeting general functions rather than specific domains, thus limiting practicality. However, BioCoder [28], a benchmark for bioinformatics code generation, has been introduced. BioCoder carefully selects 28 high-quality projects related to bioinformatics, then parses these repositories to generate ASTs and obtain relevant data. A custom code context is written for each filtered function, including necessary imports, cross-file dependencies, and some test cases.
@@
However, all the benchmarks mentioned above have a common flaw: they all overlook the evolution of projects over time. These benchmarks take the latest version of the project as the context source, and such an evolution-ignored setting makes it difficult for them to accurately reflect the performance of LLMs in real-world development scenarios. Therefore, in this work, we propose HumanEvo, the first repository-level evolution-aware code generation benchmark. Before obtaining the project context for the target function, we roll back the repository to the state before the target function was committed to ensure that the project context provided to LLMs does not suffer from issues such as future context leakage and missing useful context. Our objective is to offer LLMs a more realistic development scenario, thus more accurately reflecting the actual performance of LLMs in real-world development scenarios.
VII-B Repository-level Code Generation
Many technical studies have been proposed to better tackle repository-level code generation tasks [4, 2, 6, 48, 49, 50, 65, 66, 67, 68, 69, 77, 78]. DocPrompting [2] uses natural language (NL) intent to retrieve documents related to the target code and then sends NL intent and retrieved documents to LLMs. CoCoMIC [1] develops a cross-file context finder, which is a static code analysis tool used to retrieve the most related cross-file context. MGD [3] leverages IDEs to assist in providing context. IDE is invoked to provide useful suggestions, when reaching the pre-defined trigger point. RLPG [5] analyses dependencies potentially existing in the target function, then extracts project context that may be conducive to generating the target function. RepoCoder [6] first slices the source code files in the repository to construct a retrieval library, then uses the generated problem as a query, calculates similarity to obtain project context, and feeds the project context and problem to the model to generate initial results. After obtaining the initial results, the previous generation results are used as queries for further querying to obtain more context. This process is repeated to continuously optimize the generation results. $A^{3}$-CodGen [7] first constructs a repository knowledge base covering knowledge of third-party libraries and extracts information on mobile and code context. During the code generation process, LLM first generates an initial version of code based on NL intent, then uses NL intent and the initial version of code to retrieve as much foundational knowledge as possible from the knowledge base for the LLM. Finally, local, global, and third-party library information is integrated into the prompt to allow the model to generate the final results.
All the mentioned works treat the latest version of the project as the context source, falling into the category of evolution-ignored scenarios. Although in RepoCoder, researchers begin to take notice of this issue by considering all code after the target function in the corresponding file as “future context”, this still falls short of reality and introduces a lack of the context that should be present. Therefore, we propose a novel repository-level code generation benchmark, HumanEvo, an evolution-aware pragmatic benchmark, to better simulate real development scenarios for LLMs.
VII-C LLMs for Code Generation
CodeLlama [8] is further trained based on Llama2 [23]. It is capable of understanding and generating code in various programming languages, including but not limited to Python, Java, JavaScript, etc. CodeLlama not only generates complete code but also provides suggestions and optimization solutions for specific code snippets, greatly assisting developers. During the training of the base model, weights are initialized with an equivalent parameter amount of the Llama2 model, followed by training on a dataset of 500 billion tokens of code. StarCoder [9] is derived from StarCoderBase, which is trained using 10 trillion tokens from The Stack, a vast collection of licensed GitHub repositories with inspection tools and selection exit procedures. The DeepSeek-Coder [10] takes into account the often intricate dependency relationships inherent in real-world projects. The model needs to fully consider these dependencies before completing programming tasks. Therefore, to equip DeepSeek-Coder to handle real-world complex projects, the training data is structured to support repository-level comprehension, enhancing the model’s understanding of complex projects. This enables the model to utilize cross-file context effectively. MetaGPT [24] constructs a multi-agent collaborative framework, employing Standardized Operating Procedures and involving multiple roles such as programmers and product managers to collaborate on project completion. The results generated by agents are not saved in conversational form but in structured output formats such as charts, files, etc. Communication between agents is also conducted through structured files. MetaGPT also implements an iterative development pipeline, optimizing results through multiple rounds of iteration by recording information generated during the development process by different roles. Similarly, ChatDev [25] adopts a multi-agent framework, allocating and assigning different AI agents to various job functions. ChatDev facilitates communication among different role agents, forming a complete software development process where each agent performs its respective tasks while collaborating uniformly. In addition to these models, there are many other excellent models used in the field of code generation [29, 30, 31, 17, 32, 33, 18].
VIII Conclusion
@@
In this paper, we highlight a previously overlooked flaw in evaluating existing repository-level code generation methods: the evolution of projects over time. To address it, we introduce HumanEvo, an evolution-aware repository code generation benchmark, and conduct a comprehensive empirical study to explore LLMs’ performance in real-world development scenarios. Experimental results demonstrate that the evolution-ignored situation leads to the inflated performance of LLMs. By incorporating temporal considerations into the benchmark, HumanEvo provides a more comprehensive assessment of model performance in pragmatic development scenarios. We believe that our empirical results and HumanEvo will offer valuable insights for the future development of LLMs-based repository-level code generation methods and benchmarks.
IX Acknowledgements
This work is supported by the National Natural Science Foundation of China (62032025), the Guangdong Basic and Applied Basic Research Foundation (2023A1515012292) and CCF-Huawei Populus Grove Fund CCFHuaweiSE202403.